OpenStudy (anonymous):

A person borrows $15,000 to be paid back at 5% interest (compounded) over 60 months. Makes 6 payments of $270, then borrows an additional $40,000. Makes 7 payments of $1,030, then borrows $6,000. Makes 4 payments of $1,300. How much do they owe now and can you tell me how you came up with the answer?

OpenStudy (tkhunny):

What's your plan to solve this? What tools have we? Paper and pencil? Calculator? Spreadsheet? Super Computer? Actuary?

OpenStudy (anonymous):

paper and pencil, simple calculator, and a loan calculator. HELP!

OpenStudy (tkhunny):

Loan calculator? Set it to "Payments at the End of the Period" and we can get started.

OpenStudy (anonymous):

Rats, I was going to cut and paste but can't. Ok, borrowed $15,000 in June '11. Paid $620. Borrowed an additional $40,000 in Jan 2012. Paid $7,210. In Sept borrowed an additional $6,000. Paid $4,120. Hence, has borrowed a total of $61,000 and has made payments totalling $11,950

OpenStudy (anonymous):

Any thoughts? And the $620 should be $1,220.

OpenStudy (tkhunny):

Which one are we doing? I'm tempted to start with the first one. Set up the monthly interest factor. \(i = 0.05\) \(j = i/12 = 0.05/12 = 0.0041\overline{6}\) \(r = 1+j = 1.0041\overline{6}\)

OpenStudy (anonymous):

Ok, so the person owes 1.00416 X the balance?

OpenStudy (tkhunny):

No. Patience. We need to undertand one piece at a time. Do you understand that this is the MONTHLY interest rate factor to accumulate the balance?

OpenStudy (tkhunny):

Time - 0 Months - $15,000.00 - New Loan Time = 1 Month - $15,000.00*r - $270.00 = $14,792.50 Time = 2 Months - $14,792.50*r - $270.00 = $14,584.14 .Do this for 6 Payments.

OpenStudy (anonymous):

Thank you so much :)!

OpenStudy (tkhunny):

Well, that's just a start. These are sufficiently complicated that you may just have to step through them. A spreadsheet would be ideal.

OpenStudy (anonymous):

Ok, so when I get to 6 months, I don't just add $40,000 to the balance?

OpenStudy (tkhunny):

Do be careful with "months" vs. "payments". They will not always be the same thing. In this case, for the first six months, that will be okay. In other words, yes, after the 6th payment, add $40,000.00 to the balance. After 7 more payments, add $6,000.00 to the balance. etc.

OpenStudy (anonymous):

Correct. Thanks again so much !